Earth orbits the sun once every 365.25 days.




Find the average angular speed of Earth




about the sun.




Answer in units of rad/s.

1
Expert's answer
2022-04-28T16:42:14-0400
"\omega=\frac{2\pi}{T}=\frac{6.28}{365.25*24*3600\:\rm s}=2.0*10^{-7}\:\rm rad/s"